In this extra credit assingment you are to find the area of an equilateral triangle with each side 2 using S Sp dördr
Answer #1

y A(1,13) Los B TA O (210) equation of rine AB » 9-13=-13 (0-1) In Polar tout coordinates y=rsine; s=rcoso is equation of lin3 olr ca3 Here, OLALT Isino + 13 cos 6) lsinotrucoso r dr do 23 Is dorar T 3 s 0 0 r 3 ) 3 2. * (253) de 2 (sino + 13 (OSA) 2T 17 = 2 0 odo 6 seco do Sino+v3 050) (20n 6+13) Now Put tanot 3 = Z Gdz .. ?V3 Pv3 - 2 flere IN Z + RUB V3 V3 V3 6. 23 73Area of equilateral triangre le with each side 2 is equal to 13.